# suppSupplementary References

### v2

@@ Abstract
A/B testers that conduct large-scale tests often prioritize lifts as the main outcome metric and want to be able to control costs resulting from false rejections of the null. This work develops a decision-theoretic framework for maximizing profits subject to false discovery rate (FDR) control. We build an empirical Bayes solution for the problem via a greedy knapsack approach. We derive an oracle rule based on ranking the ratio of expected lifts and the cost of wrong rejections using the local false discovery rate (lfdr) statistic. Our oracle decision rule is valid and optimal for large-scale tests. Further, we establish asymptotic validity for the data-driven procedure and demonstrate finite-sample validity in experimental studies. We also demonstrate the merit of the proposed method over other FDR control methods. Finally, we discuss an application to data collected by experiments on the Optimizely platform.
@@ 1 Introduction
A/B testing is often used to optimize websites, marketing campaigns, and user interfaces. Leading companies report running over 10,000 experiments annually (Kohavi and Thomke, 2017). Recently, academics introduced new extensions to the statistical tool of basic hypothesis tests that fit the practicalities of A/B tests in the field. Examples include computing profit maximizing sample sizes (Feit and Berman, 2019), designing experiments when samples are limited (Azevedo et al., 2020), and controlling for peeking behavior and optional stopping (Johari et al., 2022). One concern when running many tests is multiple hypothesis testing, which results in false discoveries. Common solutions to control for false discoveries include the Benjamini-Hochberg (BH) (Benjamini and Hochberg, 1995), Sun and Cai (SC) (Sun and Cai, 2007), and the weighted false discovery rate control (BCDS) (Basu et al., 2018) methods.
Often, A/B tests compare a new intervention to a standard practice baseline used in production. For example, comparing a new recommendation algorithm to an existing one (Kim et al., 2019; Goli et al., 2024), or trying a new targeting policy compared to an existing one (Simester et al., 2020). These experiments are sometimes called champion-challenger experiments, and their uniqueness is that the experimenter has long-term observations about the average outcomes of the baseline treatment. Further, if many such experiments are being run, the cost of switching from the champion to the challenger might differ for each experiment.
In this paper, we develop a false discovery control method that focuses on lifts (i.e., a transformation of Relative Risks) while considering the available long-term information about baseline outcomes and the potential differences in the costs of implementing the winning intervention. We take a cost-benefit approach to experimentation, aiming to maximize lift, a metric practitioners are often interested in while controlling for acceptable costs that accrue from false discoveries.
@@
We use a simulation study to demonstrate the gain in performance over existing state-of-the-art FDR control methodologies. Finally, we use data from 2,766 A/B tests conducted using a large online A/B testing platform (Optimizely) to compare the performance of our method to the other state-of-the-art methods. We hope practitioners of A/B testing will find our method useful when their objective is to optimize lifts in Champion-Challenger experiments subject to cost control.
@@ 1.1 False Discovery Rate Control in A/B Testing
In large-scale experiments that test many hypotheses simultaneously, false discovery rate (FDR) control is often recommended to provide more statistical power over the family-wise error rate (FWER), which controls the probability of having at least one false positive but does not control the expected false discovery proportions. Berman and Van den Bulte (2022) reported up to 25% two-sided error proportions when practitioners do not correct for multiple testing at 5% significance. The inflation in error proportion is alarming and can lead to considerable wastage by switching to null interventions with implementation costs. Although false discoveries by themselves might not have a cost, if each test is weighted by the implementation costs, then the expected value of the ratio of false discoveries to the number of discoveries made (which is the definition of FDR) can be viewed as a proportion of ‘wasted’ investment resources. This phenomenon is more severe when the proportion of true null effects is higher. In our simulations, we find that for 80% null effects, up to 50% discoveries are false (one-sided) when not appropriately corrected at 5% significance.
Our approach to FDR control is based on a line of work by Sun and Cai (2007), Efron (2008) and Basu et al. (2018), which relies on empirical Bayes estimation and adjustments, leading to a more robust methodology that also considers variations of power as an objective function to maximize discoveries.
@@ 1.2 Relative Risk and Lift for Two-Sample t-Tests
Relative risk has been a common metric of interest in epidemiology and medicine research. It is often not recommended as the sole metric of interest but as part of an evaluation in conjunction with absolute risk or risk difference. In the context of A/B testing, relative risk surfaces through the interest of companies in comparing their experiments using lifts (which are the relative risks minus one). In this work, we study the impact of ranking multiple A/B test results by their expected lifts to decide which ones to implement. However, we anticipate our work to be helpful in epidemiological studies, economics, finance, and marketing. Irrespective of the domain of expertise, we refer to the users of A/B tests as experimenters and assume that the data has been generated from randomized trials.
@@ 1.3 Executive Summary
  • An FDR control approach, such as BH or SC (Benjamini and Hochberg, 1995; Sun and Cai, 2007) is recommended to avoid inflated false discoveries when testing multiple A/B tests. The correction is necessary even if the experiment has a second stage, replicating only the best variation in each test and following a first stage with several variations in each experiment.
  • If experimenters are interested in using lifts as the primary metric, it is recommended that they do so by applying a logarithm transformation to the relative risk to ensure asymptotic normality. Further, a second-order mean correction of half the estimated variance is necessary to de-bias the conventional estimated mean. The p-value or t-statistics from this process can then be used for FDR control.
  • To accommodate the heterogeneity in baseline profits, we derived an optimal oracle testing procedure via a greedy knapsack approach, which we call Ranking by Lifts (RBL). In our data-driven application, and in practice, the empirical performance in FDR control validity matches that of the BCDS (Basu et al., 2018) weighted multiple testing procedure while often providing a considerable performance improvement. We recommend using our proposed RBL procedure when profit considerations are important since they can potentially provide a sizeable gain in net profit over conventional FDR control methods such as standard BH, SC, or BCDS.
@@ 2 Problem Formulation
The objective of the experimenter is to select values for $\delta_{i}$ to maximize $\mathbb{E}\sum_{i}\Pi_{i}\ell_{i}\delta_{i}$, subject to the constraint $\mathbb{E}\{\sum(1-\theta_{i})\delta_{i}c_{i}/\sum\delta_{i}c_{i}\}\leq\alpha$. The value $\Pi_{i}$ indicates the profit the firm makes from sending the champion treatment to the customers, including the conversion rate and the profit margin, which is multiplied by the number of customers. Hence, switching to the challenger will add an incremental profit from the improvement in conversion rates. The expression for the constraint indicates that the expected proportional cost of erroneous switching decisions is bounded by $\alpha$.
Each experiment has sample sizes $n_{i0}$ and $n_{i1}$ and observed conversions $Y_{i0}$ and $Y_{i1}$. The experimenter computes the estimated conversion rates $\widehat{p}_{i0}=Y_{i0}/n_{i0}$ and similarly for $i1$, yielding the estimated lifts ^$\widehat{\ell}_{i}=\widehat{p}_{i1}/\widehat{p}_{i0}-1$.
@@
[…]
@@ 3.1 Deriving the Value and Weight Components
Following Equation 2.7 in Lehmann and Romano (2005) which we modify using the conventions of Efron (2011), we define a natural exponential family of distributions $f_{\zeta}(z)=C(\zeta)e^{\zeta z}f_{0}(z)$. Invoking Bayes’ rule, the posterior density of $\zeta|z$ is $g(\zeta|z)=f_{\zeta}(z)g(\zeta)/f(z)$, where $f(z)$ is the marginal density and $\zeta\sim g(\cdot)$ is a one-group model with possible atoms. We then rewrite $g(\zeta|z)=C(\zeta)g(\zeta)e^{\zeta z-\ln(f(z)/f_{0}(z))}$.
We rename $A(z):=\ln(f(z)/f_{0}(z))$, noticing that the $lfdr$ is $\pi_{0}e^{-A(z)}$. The moment generating function (MGF) of $\zeta|z$ can be written as $e^{A(z+1)-A(z)}$, which we rewrite as $lfdr(z)/lfdr(z+1)$ for convenience of estimation later. Recalling that the quantity of interest to us in the objective function is $E[e^{\zeta}-1|z]$, then $E[e^{\zeta}-1|z]=lfdr(z)/lfdr(z+1)-1$. The weight component $\mathbb{E}[c_{i}(1-\theta_{i}-\alpha)|data]$ is connected to the lfdr by the expression $c_{i}(lfdr_{i}-\alpha)$.
Figure 1 provides a graphical illustration of the value and weight components for a mixture distribution where $F_{0}$ is $N(0,1)$ and $F_{1}$ is a mixture of $N(-2,1)$ and $N(2,1)$ with equal probability. The cost function is symmetric around 0 for different values of $z$. The value function is not symmetric. For very negative values of $z$, the value is negative. However, these hypotheses would contribute a positive value to the objective function for a mildly negative value of $z$ (roughly half of the standard deviation on the left of the vertical axis at zero).
@@
The figure displays the value and cost functions for z-values ranging from [-5, 5]. The underlying data-generating process is assumed to be from a normal mixture with a 0.8 probability of being the null of N(0, 1) and non-null distributions of N(2, 1) and N(-2, 1) with equal remaining mass. We choose $\alpha=0.05$ for computing the cost. We note that roughly at $z=-4$, both the value and cost are negative; at $z=-2$, the value is still negative, and the cost becomes positive; at $z=2$, the value turns positive while the cost stays positive, and at $z=4$ the value stays positive, and the cost decreases below zero.
Remark: The definition of the lfdr we used assumes that $Z\sim\mathcal{N}(0,1)$ is symmetric under the null. In our case, some experiments have negative lifts that will generate low lfdr, although they should not be rejected (as they have a high weight). To accommodate that, we adjust the null density for $Z$ to be evaluated at $\sup_{\zeta\leq 0}N(\zeta,1)$ which is essentially the same as $N(max(z,0);0,1)$. Note that we do not use this adjustment for the ‘value’ statistics but only for the ‘weight’ statistics. [^1] The null probability $\pi_{0}=Pr(\zeta\leq 0)$ might also be estimated as $Pr(\zeta=0)+(1-Pr(\zeta=0))/2$ but, in our numeric experiments we did not find sufficient practical difference in power and validity.
@@ 3.2 The Large Sample Distribution of $Z=\ln\hat{p}_{1}/\hat{p}_{0}$
In the previous derivation, we assumed a large enough sample for $Z_{i}$ that facilitates the parametric exponential family distribution assumption of $f_{\zeta}(\cdot)$. This approach is popular in medical studies (Katz et al., 1978), and in this Section, we provide details to justify this assumption. The asymptotic distribution of $Z_{i}$ when $n_{0i},n_{1i}\to\infty$ is:
@@ 3.3 The Proposed Greedy Knapsack Algorithm
For the hypotheses with either negative weight and negative value (quadrant I) or positive weight and positive value (quadrant III), we need to decide whether to include or exclude them based on the ratio of their value to weight. We would like to include those that either provide a high positive value to positive weight ratio (thus not taking too much space but adding a lot of value) or those with a low negative value to negative weight (therefore not reducing the total value in the knapsack by much, but adding capacity to the knapsack). The order of inclusion or exclusion is done by the rank of the value-to-weight ratio, which we prove is optimal in Section 4.1. The details of the algorithm are described below.
@@
Decision Rule Weight $>$ 0 Weight $<$ 0
Value $>$ 0 II: Exclude and rank to decide I: Always reject the null
Value $<$ 0 III: Always accept the null IV: Include and rank to decide
The algorithm will a) always include in the knapsack (reject the null) hypotheses in quadrant I, b) always exclude (fail to reject) in quadrant III, and c) decide on hypotheses in quadrants II and IV according to their ranking of value to weight ratio.
Ranking by Lifts (RBL) Algorithm Step 1. For each hypothesis compute the weight $c_{i}(lfdr_{i}(h_{i})-\alpha)$, and the value $\Pi_{i}(lfdr_{i}(h_{i})/lfdr_{i}(h_{i}+\hat{\sigma_{i}})-1)$. Initially, the knapsack is empty with capacity zero. First, all hypotheses with negative weight are added to the knapsack. This will enlarge the knapsack and give it a maximum capacity of $\sum_{i:lfdr_{i}<\alpha}c_{i}(\alpha-lfdr_{i}).$ Step 2. Discard all hypotheses with positive weight and negative value from consideration, as they will take up knapsack space and reduce the total value of items in the knapsack. Step 3. Create a consideration set of the items with either negative weight and value or positive weight and value, and rank them based on the value-to-weight ratio $\frac{\Pi_{i}(lfdr_{i}(h_{i})/lfdr_{i}(h_{i}+\hat{\sigma}_{i})-1)}{c_{i}(lfdr_{i}(h_{i})-\alpha)}.$ Step 4. For the items in the consideration set, start with the highest ranked item, and for each item: (i) If it has a negative weight and value (which means it was added in Step 1), discard it from the knapsack. (ii) Add it if it has positive weight and value (which means it is currently outside the knapsack). Stop when there is no more knapsack capacity available.

### v3

@@ Abstract
A/B testing is a core tool for decision-making in business experimentation, particularly in digital platforms and marketplaces. Practitioners often prioritize lift in performance metrics while seeking to control the costs of false discoveries. This paper develops a decision-theoretic framework for maximizing expected profit subject to a constraint on the cost-weighted false discovery rate (FDR). We propose an empirical Bayes approach that uses a greedy knapsack algorithm to rank experiments based on the ratio of expected lift to cost, incorporating the local false discovery rate (lfdr) as a key statistic. The resulting oracle rule is valid and rank-optimal. In large-scale settings, we establish the asymptotic validity of a data-driven implementation and demonstrate superior finite-sample performance over existing FDR-controlling methods. An application to A/B tests run on the Optimizely platform highlights the business value of the approach.
@@ 1 Introduction
A/B testing is often used to optimize websites, marketing campaigns, and user interfaces. Leading companies report running over 10,000 experiments annually (Kohavi and Thomke, , 2017). Recently, academics introduced new extensions to the statistical tool of basic hypothesis tests that fits the practicalities of A/B tests in the field. Examples include computing profit-maximizing sample sizes (Feit and Berman, , 2019), designing experiments when samples are limited (Azevedo et al., , 2020), and controlling for peeking behavior and optional stopping (Johari et al., , 2022). One concern when running many tests is multiple hypothesis testing, which results in false discoveries. Common solutions to control for false discoveries include the Benjamini-Hochberg (BH) (Benjamini and Hochberg, , 1995), Sun and Cai (SC) (Sun and Cai, , 2007), and the weighted false discovery rate control (BCDS) (Basu et al., , 2018) methods.
We base our framework on a prevalent setup in the literature (Azevedo et al., , 2020; Sudijono et al., , 2024) where a collection of ideas, often having a common connection, are tested using limited resources. The authors in the latter work call the setup ‘experimentation programs’. In these programs, many A/B tests are run concurrently, and each test often compares a new intervention to a standard practice baseline used in production. For example, comparing a new recommendation algorithm to an existing one (Kim et al., , 2019; Goli et al., , 2024), or trying a new targeting policy compared to an existing one (Simester et al., , 2020). These experiments are sometimes called champion-challenger experiments, and their uniqueness is that the experimenter has long-term observations about the average outcomes of the baseline treatment. Further, if many such experiments are being run, the cost of switching from the champion to the challenger might differ for each experiment.
In this paper, we develop a false discovery control method that focuses on lifts (i.e., a transformation of Relative Risks) while considering the available long-term information about baseline outcomes and the potential differences in the costs of implementing the winning intervention. We take a cost-benefit approach to experimentation, aiming to maximize lift, a metric practitioners are often interested in, while controlling for acceptable costs that accrue from false discoveries.
@@
We use a simulation study to demonstrate the gain in performance over existing state-of-the-art FDR control methodologies. Finally, we use data from 2,766 A/B tests conducted using a large online A/B testing platform (Optimizely) to compare the performance of our method to the other state-of-the-art methods. We hope practitioners of A/B testing will find our method useful when their objective is to optimize lifts in Champion-Challenger experiments, subject to cost control.
@@ 1.1 False Discovery Rate Control in A/B Testing
In large-scale experiments that test many hypotheses simultaneously, false discovery rate (FDR) control is often recommended to provide more statistical power over the family-wise error rate (FWER), which controls the probability of having at least one false positive but does not control the expected false discovery proportions. Berman and Van den Bulte, (2022) reported up to 25% two-sided error proportions when practitioners do not correct for multiple testing at 5% significance. The inflation in error proportion is alarming and can lead to considerable wastage by switching to null interventions with implementation costs. Although false discoveries by themselves might not have a cost, if each test is weighted by the implementation costs, then the expected value of the ratio of false discoveries to the number of discoveries made (which is the definition of FDR) can be viewed as a proportion of ‘wasted’ investment resources. This phenomenon is more severe when the proportion of true null effects is higher. In our simulations, we find that for 80% null effects, up to 50% discoveries are false (one-sided) when not appropriately corrected at 5% significance.
Our approach to FDR control is based on a line of work by Sun and Cai, (2007), Efron, (2008), and Basu et al., (2018), which relies on empirical Bayes estimation and adjustments, leading to a more robust methodology that also considers variations of power as an objective function to maximize discoveries.
@@ 1.2 Relative Risk and Lift for Two-Sample t-Tests
Relative risk has been a common metric of interest in epidemiology and medical research. It is often not recommended as the sole metric of interest but as part of an evaluation in conjunction with absolute risk or risk difference. In the context of A/B testing, relative risk surfaces through the interest of companies in comparing their experiments using lifts (which are the relative risks minus one). In this work, we study the impact of ranking multiple A/B test results by their expected lifts to decide which ones to implement. However, we anticipate our work to be helpful in epidemiological studies, economics, finance, and marketing. Irrespective of the domain of expertise, we refer to the users of A/B tests as experimenters and assume that the data has been generated from randomized trials.
@@ 1.3 Executive Summary
  • An FDR control approach, such as BH or SC (Benjamini and Hochberg, , 1995; Sun and Cai, , 2007) is recommended to avoid inflated false discoveries when testing multiple A/B tests. The correction is necessary even if the experiment has a second stage, replicating only the best variation in each test and following a first stage with several variations in each experiment.
  • If experimenters are interested in using lifts as the primary metric, it is recommended that they do so by applying a logarithm transformation to the relative risk to ensure asymptotic normality. Further, a second-order mean correction of half the estimated variance is necessary to de-bias the conventional estimated mean. The p-value or t-statistic from this process can then be used for FDR control.
  • To accommodate the heterogeneity in baseline profits, we derived an optimal oracle testing procedure via a greedy knapsack approach, which we call Ranking by Lifts (RBL). In our data-driven application, and in practice, the empirical performance in FDR control validity matches that of the BCDS (Basu et al., , 2018) weighted multiple testing procedure while often providing a considerable power improvement. We recommend using our proposed RBL procedure when profit considerations are important since they can potentially provide a sizeable gain in net profit over conventional FDR control methods such as standard BH, SC, or BCDS.
@@ 2 Problem Formulation
The objective of the experimenter is to select values for $\delta_{i}$ to maximize $\mathbb{E}\sum_{i}\Pi_{i}\ell_{i}\delta_{i}$, subject to the constraint $\mathbb{E}\{\sum(1-\theta_{i})\delta_{i}c_{i}/\sum\delta_{i}c_{i}\}\leq\alpha$. The value $\Pi_{i}$ indicates the profit the firm makes from sending the champion treatment to the customers, including the conversion rate and the profit margin, which is multiplied by the number of customers. Hence, switching to the challenger will add an incremental profit from the improvement in conversion rates. The expression for the constraint indicates that the expected proportional cost of erroneous switching decisions is bounded by $\alpha$. This formulation allows the experimenter to trade off longer-term learning about interventions that have true effects, with maximizing short-term net-profit that can reject too many or too few hypotheses. [^1] Appendix A analyzes an alternate net-profit objective and provides more clarity behind the rationale for our objective of choice. Each experiment has sample sizes $n_{i0}$ and $n_{i1}$ and observed conversions $Y_{i0}$ and $Y_{i1}$. The experimenter computes the estimated conversion rates $\widehat{p}_{i0}=Y_{i0}/n_{i0}$ and similarly for $i1$, yielding the estimated lifts $\widehat{\ell}_{i}=\widehat{p}_{i1}/\widehat{p}_{i0}-1$.
@@
$\max_{\delta_{i}\in\{0,1\}}\mathbb{E}\sum_{i}\Pi_{i}\ell_{i}\delta_{i}\text{~~subject to~~}\mathbb{E}\{\sum(1-\theta_{i})\delta_{i}c_{i}/\sum\delta_{i}c_{i}\}\leq\alpha.$
@@ 3.1 Deriving the Value and Weight Components
Following Equation 2.7 in Lehmann and Romano, (2005) which we modify using the conventions of Efron, (2011), we define a natural exponential family of distributions $f_{\zeta}(z)=C(\zeta)e^{\zeta z}f_{0}(z)$. Invoking Bayes’ rule, the posterior density of $\zeta|z$ is $g(\zeta|z)=f_{\zeta}(z)g(\zeta)/f(z)$, where $f(z)$ is the marginal density and $\zeta\sim g(\cdot)$ is a one-group model with possible atoms. We then rewrite $g(\zeta|z)=C(\zeta)g(\zeta)e^{\zeta z-\ln(f(z)/f_{0}(z))}$.
We rename $A(z):=\ln(f(z)/f_{0}(z))$, noticing that the lfdrlfdr is $\pi_{0}e^{-A(z)}$. The moment generating function (MGF) of $\zeta|z$ can be written as $e^{A(z+1)-A(z)}$, which we rewrite as $lfdr(z)/lfdr(z+1)$ for convenience of estimation later. Recalling that the quantity of interest to us in the objective function is $E[e^{\zeta}-1|z]$, then $E[e^{\zeta}-1|z]=lfdr(z)/lfdr(z+1)-1$. The weight component $\mathbb{E}[c_{i}(1-\theta_{i}-\alpha)|data]$ is connected to the lfdr by the expression $c_{i}(lfdr_{i}-\alpha)$.
Figure 1 provides a graphical illustration of the value and weight components for a mixture distribution where $F_{0}$ is $N(0,1)$ and $F_{1}$ is a mixture of N(2,1)N(-2,1) and $N(2,1)$ with equal probability. The cost function is symmetric around 0 for different values of $z$. The value function is not symmetric. For very negative values of $z$, the value is negative. However, these hypotheses would contribute a positive value to the objective function for a mildly negative value of $z$ (roughly half of the standard deviation on the left of the vertical axis at zero).
@@
The figure displays the value and cost functions for z-values ranging from [-5, 5]. The underlying data-generating process is assumed to be from a normal mixture with a 0.8 probability of being the null of N(0, 1) and non-null distributions of N(2, 1) and N(-2, 1) with equal remaining mass. We choose $\alpha=0.05$ for computing the cost. We note that roughly at z=4z=-4, both the value and cost are negative; at z=2z=-2, the value is still negative, and the cost becomes positive; at $z=2$, the value turns positive while the cost stays positive, and at $z=4$ the value stays positive, and the cost decreases below zero.
Remark: The definition of the lfdr we used assumes that $Z\sim\mathcal{N}(0,1)$ is symmetric under the null. In our case, some experiments have negative lifts that will generate low lfdr, although they should not be rejected (as they have a high weight). To accommodate that, we adjust the null density for $Z$ to be evaluated at $\sup_{\zeta\leq 0}N(\zeta,1)$ which is essentially the same as $N(max(z,0);0,1)$. Note that we do not use this adjustment for the ‘value’ statistics but only for the ‘weight’ statistics. [^2] The null probability $\pi_{0}=Pr(\zeta\leq 0)$ might also be estimated as $Pr(\zeta=0)+(1-Pr(\zeta=0))/2$ but, in our numeric experiments, we did not find sufficient practical difference in power and validity.
@@ 3.2 The Large Sample Distribution of $Z=\ln\hat{p}_{1}/\hat{p}_{0}$
In the previous derivation, we assumed a large enough sample for $Z_{i}$ that facilitates the parametric exponential family distribution assumption of $f_{\zeta}(\cdot)$. This approach is popular in medical studies (Katz et al., , 1978), and in this Section, we provide details to justify this assumption. The asymptotic distribution of $Z_{i}$ when $n_{0i},n_{1i}\to\infty$ is:
@@ 3.3 The Proposed Greedy Knapsack Algorithm
For the hypotheses with either negative weight and negative value (quadrant IV) or positive weight and positive value (quadrant II), we need to decide whether to include or exclude them based on the ratio of their value to weight. We would like to include those that either provide a high positive value to positive weight ratio (thus not taking too much space but adding a lot of value) or those with a low negative value to negative weight (therefore not reducing the total value in the knapsack by much, but adding capacity to the knapsack). The order of inclusion or exclusion is done by the rank of the value-to-weight ratio, which we prove is optimal in Section 4.1. The details of the algorithm are described below.
@@
Decision Rule Weight >> 0 Weight << 0
Value >> 0 II: Exclude and rank to decide I: Always reject the null
Value << 0 III: Always accept the null IV: Include and rank to decide
The algorithm will a) always include in the knapsack (reject the null) hypotheses in quadrant I, b) always exclude (fail to reject) in quadrant III, and c) decide on hypotheses in quadrants II and IV according to their ranking of value-to-weight ratio.
Ranking by Lifts (RBL) Algorithm Step 1. For each hypothesis compute the weight $c_{i}(lfdr_{i}(h_{i})-\alpha)$, and the value $\Pi_{i}(lfdr_{i}(h_{i})/lfdr_{i}(h_{i}+\hat{\sigma_{i}})-1)$. Initially, the knapsack is empty with a capacity of zero. First, all hypotheses with negative weight are added to the knapsack. This will enlarge the knapsack and give it a maximum capacity of $\sum_{i:lfdr_{i}<\alpha}c_{i}(\alpha-lfdr_{i}).$ Step 2. Discard all hypotheses with positive weight and negative value from consideration, as they will take up knapsack space and reduce the total value of items in the knapsack. Step 3. Create a consideration set of the items with either negative weight and value or positive weight and value, and rank them based on the value-to-weight ratio $\frac{\Pi_{i}(lfdr_{i}(h_{i})/lfdr_{i}(h_{i}+\hat{\sigma}_{i})-1)}{c_{i}(lfdr_{i}(h_{i})-\alpha)}.$ Step 4. For the items in the consideration set, start with the highest-ranked item, and for each item: (i) If it has a negative weight and value (which means it was added in Step 1), discard it from the knapsack. (ii) Add it if it has positive weight and value (which means it is currently outside the knapsack). Stop when there is no more knapsack capacity available.